OpenStudy (anonymous):

If A be a point on the ellipse x^2/a^2 + y^2/b^2 = 1 and B be a point on its auxillary circle vertically above it, then find the locus of the midpt of A,B

OpenStudy (anonymous):

You talk like a pirate :) Ye be saying "be" instead of "is" me hardy!

OpenStudy (anonymous):

\[A=(a \cos \theta ,bsin \theta ) \]\\\[ B=(acos \theta , asin \theta )\] midpoint p=\[(acos \theta , (a+b)\sin \theta /2)\] therefore (x/a)^2+(2y/(a+b))^2=1. that is the locus........which is again a ellipse,

OpenStudy (anonymous):

gr8 man

OpenStudy (anonymous):

the other ones